OpenStudy (anonymous):

Find the inverse of the function. f(x) = 4x3 - 7

OpenStudy (sparrow2):

\[f-1()=\sqrt[s]{(x+7)/}\]

OpenStudy (sparrow2):

cube root

OpenStudy (anonymous):

Not a one-to-one function f-1(x) = cube root of x minus divided by four plus seven f-1(x) = cube root of quantity x plus seven divided by four f-1(x) = cube root of quantity x minus seven divided by four

OpenStudy (sparrow2):

oh devided by 4

OpenStudy (sparrow2):

3

OpenStudy (anonymous):

so its f-1(x) = cube root of quantity x plus seven divided by four

OpenStudy (anonymous):

@sparrow2

OpenStudy (sparrow2):

yes